# Pre-Trained Vision-Language Model Selection and Reuse for Downstream Tasks

### v1

@@ 1 Introduction
Vision-Language Models (VLMs), such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), etc, which are pre-trained on large-scale image-text datasets, have recently attracted significant attention due to their remarkable zero-shot prediction capabilities on visual tasks. However, though VLM shows impressive general ability, as highlighted in Radford et al. (2021), VLMs often fall short of supervised expert models in many downstream tasks. To address this limitation, numerous studies (Dosovitskiy et al., 2021; Yu et al., 2022; Fang et al., 2023) have sought to enhance the zero-shot performance of VLMs by studying model architectures, pre-training datasets, and training/fine-tuning methods. This effort has led to the development of many open-source pre-trained VLMs with diverse structures and parameters, contributing to VLM model hubs like open-clip (Ilharco et al., 2021), which currently hosts more than 100 pre-trained VLMs.
As more and more VLMs are open-sourced, the problem of how to select a VLM to reuse for specific downstream tasks naturally occurs. Although we can directly utilize the best-performing model on a universal dataset such as ImageNet, previous work (Fang et al., 2022) has shown that the performance of VLMs can vary greatly depending on dataset domain. For example, we evaluate the performance of various pre-trained VLMs in the open-clip library across several downstream tasks (1(a)) and within different classes of a specific task (1(b)). Figure 1(a) reveals that each VLM demonstrates distinct strengths in zero-shot visual tasks, with no single model outperforming all others across every task. Interestingly, models that perform worse on general tasks can sometimes surpass stronger models in specific downstream tasks. Furthermore, even in the same task, different VLMs exhibit varying levels of performance across specific classes, as illustrated in Figure 1(b).
@@
(b) F1 score of VLMs per class in a downstream task FER2013.
@@
Therefore, it is important to design VLM selection methods, and it would be better if we could achieve more fine-grained selection, i.e., select different VLMs to handle different classes. The direct way to select a model is to evaluate all candidate models’ performance on the target task. However, it is unrealistic due to time and computational resource limitations. Additionally, previous works on model selection (Tran et al., 2019; You et al., 2021) primarily focus on single-modal models, making them unsuitable for VLM selection since they only handle either image or text output and cannot incorporate data from the other modality. Zohar et al. (2023) is the first study to focus on VLM selection, proposing to evaluate VLM performance using textual information. However, their selection strategy heavily depends on the models’ ground-truth performance on large-scale datasets, such as ImageNet. When models excel on large-scale datasets but under-perform on specific tasks, selection strategy effectiveness drops, as shown in our experiments.
@@
To this end, we introduce a novel paradigm to select and reuse VLMs called Model Label Learning (MLL). The core idea is to organize candidate pre-trained VLMs into a model hub and describe the specialty and utility of each VLM as the model’s label in some manner. When facing a new downstream task, we can match the task requirements with the model labels to select and reuse models. Specifically, the proposal contains three key interconnected modules: model labeling, model selection, and model reuse. In the model labeling process, we construct a semantic graph with commonly occurring visual concepts and representative samples, and each model undergoes pre-testing on the semantic graph to generate its model label, which describes its capability on these semantic classes. In the model selection process, we generate caption descriptions for both the nodes in the semantic graph and the categories to be classified in the target task to compare their similarity. This enables us to evaluate the model’s performance on the target classes by aligning the matched semantic nodes with the model labels. In the model reuse process, we apply an ensemble strategy that combines the selected models’ predictions on a single class and chooses the highest confidence across all classes as the final prediction.
@@
The model labeling process is completed immediately when the candidate VLM is added to the model hub, therefore, it is target task independent, which means the proposal is both data and computationally efficient in the model selection process. Moreover, the proposal is highly growable since the capability could grow with the number of candidate models in the model hub and the model labels are also scalable since more semantic nodes can be added continually. To facilitate related research, we further introduce a comprehensive benchmark for evaluating VLM selection methods. The benchmark includes 49 pre-trained VLMs and 17 target datasets as downstream tasks. The ground-truth model ranking for each target task is provided for evaluation. We construct a semantic graph that contains more than 9000 commonly used visual concepts to pre-test VLMs. The experiments demonstrate the effectiveness of our approach in both selecting and reusing VLMs, while also validating the scalability of the model hub.
@@ 2 Related Work
Learnware (Zhou & Tan, 2022) is a novel paradigm that explores more effective model selection by constructing specifications to describe the capabilities of the model, closely aligning with our idea of model labeling. Compared with previous selection methods, learnware enables scalable and efficient model selection across diverse architectures and input types within a unified framework, improving as the system expands. Model specification is central to the learnware paradigm. Recent works (Tan et al., 2024) on learnware paradigm are built on Reduced Kernel Mean Embedding (RKME) (Wu et al., 2021), which maps training data distributions to points in Reproducing Kernel Hilbert Space (RKHS) and identifies models by comparing similarities in the RKHS. Furthermore, Guo et al. (2023) enhanced RKME for heterogeneous label spaces, while Tan et al. (2023) addressed challenges in heterogeneous feature spaces. However, learnware requires training data to construct specifications. Considering the scale of VLM pre-trained datasets, it is unrealistic to construct specifications for learnware to select models due to limited time and computational resources.
@@ 3.1 Zero-Shot Vision Task of VLM
$\hat{y}=\mathop{\arg\max}_{y\in Y_{T}}\frac{\exp\left(\operatorname{sim}\left(\mathcal{I}(x),\mathcal{T}(y)\right)\right)}{\sum\limits_{y^{\prime}\in Y_{T}}\exp\left(\operatorname{sim}\left(\mathcal{I}(x),\mathcal{T}(y^{\prime})\right)\right)}$ (1)
where $\operatorname{sim}\left(\cdot,\cdot\right)$ denotes cosine similarity.
@@ 3.2 Problem Setup
Assume the model hub has $\mathcal{M}$ pre-trained VLMs $\left\{f_{m}=\{\mathcal{I}_{m},\mathcal{T}_{m}\}\right\}_{m=1}^{\mathcal{M}}$, where $\mathcal{I}_{m}$ denote image encoder of the VLM $f_{m}$ and $\mathcal{T}_{m}$ denote text encoder of the VLM $f_{m}$. There are two stages in our setting: submission stage for model developers to upload models and identification stage for users to select models.
In submission stage, model developer submits a VLM $f_{m}$ to the model hub, and the model hub assigns a label $S_{m}$ to the model to describe its specialty and utility. It is particularly emphasized that uploaded models are anonymous, meaning we do not have access to their training data.
In identification stage, user attempts to select VLMs from the model hub for the zero-shot downstream task $T$, by uploading general information about the task, such as classes, domain type, and task type, to describe their requirements. We subsequently utilize this information to select and reuse suitable VLMs, based on the model labels established in the submission stage.
The two main problems in our settings are:
@@ 4.1 Model Labeling
$s_{m,x}^{v}=\mathrm{sim}(\mathcal{I}_{m}(x),\mathcal{T}_{m}(D_{\mathcal{G}})),x\in X_{v}$ (2)
$s_{m}^{v}=\{s_{m,x}^{v}|\;\forall x\in X_{v}\}$ (3)
@@
where $\mathcal{I}_{m}(\cdot)$, $\mathcal{T}_{m}(\cdot)$ denotes the image encoder and text encoder of model $f_{m}$
@@ 4.2 Model Selection
By utilizing the transfer matrix $Z$, the precision prediction $p_{m,y}$ for each class $y$ in the downstream task $T$ can be further derived.
@@ 5.1 MLL Benchmark
To evaluate the capabilities of the MLL paradigm in zero-shot visual tasks with VLMs, we need to obtain a set of sampling datasets for constructing semantic graph $\mathcal{G}$, along with another set dedicated to downstream target tasks. For this study, we select 49 VLMs, 5 Sample Datasets, and 17 Target Datasets. Additionally, we collect general information about task types and domains associated with each dataset to provide a task description. For testing selected models on target tasks, we utilized same prompting strategy outlined in Radford et al. (2021), ensuring consistency in our evaluation methodology. Code of MLL Benchmark and our proposal is available at this anonymous link.
@@
We utilized 5 datasets, ImageNet (Deng et al., 2009), ImageNet-V2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019), ImageNet-A (Hendrycks et al., 2021b) and ImageNet-R (Hendrycks et al., 2021a), as Sample Datasets for semantic graph construction. Additionally, we used 17 commonly used datasets and their task general information as Target Datasets to evaluate VLM selection and reuse methods in zero-shot visual tasks (as shown in  Table 5). These datasets demonstrate diversity in terms of domain, number of classes, and task types. They encompass various domains, including animals, food, text, landscapes, remote sensing, medical, and transportation. Additionally, they cover a range of tasks such as image classification, geo-localization, optical character recognition, facial expression recognition, and object distance estimation. To eliminate interference from additional modules during evaluation, all tasks can be assessed using the same VLM architecture.
@@ 5.2 Experiment Setup
We construct a semantic graph $\mathcal{G}$ containing 9055 nodes using the WordNet synsets, which contains a wide range of items, such as animals, tools, clothing, vehicles, plants, and more. Each node is represented by up to 75 randomly selected images from the sample datasets, reflecting the distribution of the node’s concepts. We use OpenAI text-embedding-3-large model to obtain caption embeddings of semantic graph nodes and downstream task class nodes, then match the similar node between them by cosine similarity between the embeddings.
@@ 5.3 Experiment Results
$k$ 1 2 3 4 5 6
Avg. 0.6620 0.6637 0.6639 0.6571 0.6556 0.6594
Interface Time Cost 1.000 1.927 2.233 2.467 2.700 2.900
@@ 6 Conclusion
In this paper, we explore how to select and reuse pre-trained VLMs for a specific downstream task. To the best of our knowledge, this problem has been rarely studied. To address this, we propose a novel paradigm called Model Label Learning (MLL) that assigns each VLM a label to describe its utility on representative visual concepts. The MLL paradigm contains three key modules: model labeling, model selection, and model reuse. The proposal is highly efficient, scalable, and convenient for both model developers and users. Moreover, we introduced a benchmark for evaluating pre-trained VLM selection and reuse methods that contain 49 pre-trained VLMs and 17 target datasets. Experiments demonstrate the proposal can achieve state-of-the-art model selection performance for VLMs and the ability to deal with downstream tasks could grow with the scale of the model hub, showing the potential of building large model hubs with advanced model selection mechanisms.

### v2

# Vision-Language Model Selection and Reuse for Downstream Adaptation

@@ 1 Introduction
Vision-Language Models (VLMs), such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), which are pre-trained on large-scale image-text datasets, have recently attracted significant attention due to their remarkable zero-shot prediction capabilities on visual tasks. However, though VLM shows impressive general ability, as highlighted in Radford et al. (2021), VLMs often fall short of supervised expert models in many downstream tasks. To address this limitation, numerous studies (Dosovitskiy et al., 2021; Yu et al., 2022; Fang et al., 2023) enhanced the zero-shot performance of VLMs by studying model architectures, pre-training datasets, and training/fine-tuning methods. This effort has led to the development of many open-source pre-trained VLMs with diverse structures and parameters, contributing to VLM model hubs like open-clip (Ilharco et al., 2021), which currently hosts more than 100 pre-trained VLMs.
With the increasing availability of open-source VLMs, selecting and reusing a pre-trained model has become a practical alternative to training one from scratch, which demands extensive training data and computing resources (Li et al., 2021), struggling with catastrophic forgetting during the incremental refinement of the trained model in open environments (Guo et al., 2025). Consequently, a critical problem naturally arises: how to select a VLM to reuse for specific downstream tasks. Although we can directly utilize the best-performing model on a universal dataset such as ImageNet, previous work (Fang et al., 2022) has shown that the performance of VLMs can vary greatly depending on dataset domain. For example, we evaluate the performance of various pre-trained VLMs in the open-clip library across several downstream tasks (1(a)) and within different classes of a specific task (1(b)). Figure 1(a) reveals that each VLM demonstrates distinct strengths in zero-shot visual tasks, with no single model outperforming all others across every task. Interestingly, models that perform worse on general tasks can sometimes surpass stronger models in specific downstream tasks. Furthermore, even in the same task, different VLMs exhibit varying levels of performance across specific classes, as illustrated in Figure 1(b).
@@
(b) F1 score of VLMs per class in FER2013.
@@
Therefore, it is important to design VLM selection methods, and it would be better if we could achieve more fine-grained selection, i.e., select different VLMs to handle different classes. The direct way to select a model is to evaluate all candidate models’ performance on the target task. However, it is unrealistic due to time and computational resource limitations. Additionally, previous works on model selection (Tran et al., 2019; You et al., 2021) primarily focus on single-modal models, making them unsuitable for VLM selection since they only handle either image or text output and cannot incorporate data from the other modality. Zohar et al. (2023) is the first study to focus on VLM selection, proposing to evaluate VLM performance using textual information. However, the selection strategy heavily depends on the models’ ground-truth performance on large-scale datasets, such as ImageNet. When models excel on large-scale datasets but under-perform on specific tasks, selection strategy effectiveness drops, as shown in our experiments.
@@
To this end, we introduce a novel paradigm to select and reuse VLMs called Model Label Learning (MLL). The core idea is to organize candidate pre-trained VLMs into a model hub and describe the specialty and utility of each VLM as the model’s label in some manner. When faced with a new downstream task, we can match the task requirements with the model labels to select and reuse models. Specifically, the proposal contains three key interconnected modules: model labeling, model selection, and model reuse. In the model labeling process, we construct a semantic graph with commonly occurring visual concepts and representative samples, and each model undergoes pre-testing on the semantic graph to generate its model label, which describes its capability on these semantic classes. In the model selection process, we generate caption descriptions for both the nodes in the semantic graph and the categories to be classified in the target task to compare their similarity. This enables us to evaluate the model’s performance on the target classes by aligning the matched semantic nodes with the model labels. In the model reuse process, we apply an ensemble strategy that combines the predictions of the selected models in a single class and chooses the highest confidence in all classes as the final prediction on the target task.
@@
The model labeling process is completed immediately when the candidate VLM is added to the model hub, therefore, it is target task independent, which means the proposal is both data and computationally efficient in the model selection process. Moreover, the proposal is highly growable since the capability could grow with the number of candidate models in the model hub and the model labels are also scalable since more semantic nodes can be added continually. To facilitate related research, we further introduce a comprehensive benchmark for evaluating VLM selection methods. The benchmark includes 49 pre-trained VLMs and 17 target datasets as downstream tasks. The ground-truth model ranking for each target task is provided for evaluation. We construct a semantic graph that contains more than 9000 commonly used visual concepts to pre-test VLMs. The experiments demonstrate the effectiveness of our approach in both selecting and reusing VLMs, while also validating the scalability of the model hub based on our proposal.
@@ 2 Related Work
Learnware (Zhou & Tan, 2022) is a novel paradigm that explores more effective model selection by constructing specifications to describe the capabilities of the model, closely aligning with our idea of model labeling. Compared with previous selection methods, learnware enables scalable and efficient model selection across diverse architectures and input types within a unified framework, improving as the system expands. Model specification is central to the learnware paradigm, describing the functionality of each model. Recent works (Tan et al., 2024) on learnware paradigm are built on Reduced Kernel Mean Embedding (RKME) (Wu et al., 2021), which maps training data distributions to points in Reproducing Kernel Hilbert Space (RKHS) and identifies models by calculating the RKHS distance between RKME specifications. Furthermore, Guo et al. (2023) enhanced RKME for heterogeneous label spaces, while Tan et al. (2023) addressed challenges in heterogeneous feature spaces. Besides, Zhou et al. (2023) extended the learnware paradigm to the domain of the diffusion models selection. However, learnware requires training data to construct specifications. Considering the scale of VLM pre-trained datasets, it is unrealistic to construct specifications for learnware to select models due to limited time and computational resources.
@@ 3.1 Zero-Shot Vision Task of VLM
$\hat{y}=\mathop{\arg\max}_{y\in Y_{T}}\frac{\exp\left(\operatorname{sim}\left(\mathcal{I}(x),\mathcal{T}(y)\right)/\tau\right)}{\sum\limits_{y^{\prime}\in Y_{T}}\exp\left(\operatorname{sim}\left(\mathcal{I}(x),\mathcal{T}(y^{\prime})\right)/\tau\right)}$ (1)
where $\operatorname{sim}\left(\cdot,\cdot\right)$ denotes cosine similarity, $\tau$ represents the temperature determined by VLMs.
@@ 3.2 Problem Setup
Assume the model hub has $\mathcal{M}$ pre-trained VLMs $\left\{f_{m}=\{\mathcal{I}_{m},\mathcal{T}_{m}\}\right\}_{m=1}^{\mathcal{M}}$, where $\mathcal{I}_{m}$ denotes the image encoder of the VLM $f_{m}$ and $\mathcal{T}_{m}$ denotes the text encoder of the VLM $f_{m}$. There are two stages in our setting: submission stage for model developers to upload models to the model hub and identification stage for users to select suitable models for their tasks from the model hub.
In submission stage, model developers submit a VLM $f_{m}$ to the model hub, and the model hub assigns a label $S_{m}$ to the model to describe its specialty and utility. It is particularly emphasized that uploaded models are anonymous, meaning we do not have access to their training data.
In identification stage, user attempts to select VLMs from the model hub for the zero-shot downstream task $T$, by uploading general information about the task, such as classes, domain type, and task type, to describe their requirements. We subsequently utilize this information to select and reuse suitable VLMs, based on the model labels which have been established in the submission stage.
Two main problems in our settings are:
@@ 4.1 Model Labeling
$s_{m,x}^{v}=\mathrm{sim}(\mathcal{I}_{m}(x),\mathcal{T}_{m}(d_{v})),x\in X_{v}$ (2)
$s_{m}^{v}=\{s_{m,x}^{v}|\;x\in X_{v}\}$ (3)
@@
where $\mathcal{I}_{m}(\cdot)$, $\mathcal{T}_{m}(\cdot)$ denotes the image encoder and the text encoder of model $f_{m}$
@@ 4.2 Model Selection
By leveraging the transfer matrix $Z$, we can further derive the precision prediction $p_{m,y}$ for each class $y$ within the downstream task $T$ as described below.
@@ 5.1 MLL Benchmark
To evaluate the capabilities of the MLL paradigm in zero-shot visual tasks with VLMs, we need to obtain a set of sampling datasets for constructing semantic graph $\mathcal{G}$, along with another set dedicated to downstream target tasks. For this study, we select 49 VLMs, 5 Sample Datasets, and 17 Target Datasets. Additionally, we collect general information about task types and domains associated with each dataset to provide a task description. For testing selected models on target tasks, we utilized same prompting strategy outlined in Radford et al. (2021), ensuring consistency in our evaluation methodology.
@@
We utilized 5 datasets, ImageNet (Deng et al., 2009), ImageNet-V2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019), ImageNet-A (Hendrycks et al., 2021b) and ImageNet-R (Hendrycks et al., 2021a), as Sample Datasets for semantic graph construction. Additionally, we used 17 commonly used datasets and their task general information as Target Datasets to evaluate VLM selection and reuse methods in zero-shot visual tasks (as shown in  Table 5). These datasets demonstrate diversity in terms of domain, number of classes, and task types. They encompass various domains, including animals, food, text, landscapes, remote sensing, medical, and transportation. Additionally, they cover a range of tasks such as image classification, geo-localization, optical character recognition, facial expression recognition, land cover classification, and object distance estimation. To eliminate interference of predication from additional modules during evaluation, all tasks can be assessed using the same VLM architecture.
@@ 5.2 Experiment Setup
In order to build a rich semantic graph $\mathcal{G}$, we used WordNet synsets to carefully select 9055 nodes, covering many fields such as animals, tools, clothing, vehicles, plants, ensuring the representativeness of the semantic graph. We randomly selected up to 75 pictures for each node, which are all from the sample datasets and reflected the distribution of the node’s concepts. In addition, we used the OpenAI text-embedding-3-large model to obtain their caption embeddings. By calculating the cosine similarity between these embeddings, we can accurately match the correspondence between the semantic graph nodes and the downstream task classes.
@@ 5.3 Experiment Results
$k$ Average Accuracy Interface Time Cost
@@ 6 Conclusion
In this paper, we explore how to select and reuse pre-trained VLMs for a specific downstream task. To the best of our knowledge, this problem has rarely been studied. To address this, we propose a novel paradigm called Model Label Learning (MLL) that assigns each VLM a label to describe its utility on representative visual concepts. The MLL paradigm contains three key modules: model labeling, model selection, and model reuse. The proposal is highly efficient, scalable, and convenient for both model developers and users. Moreover, we introduced a benchmark for evaluating pre-trained VLM selection and reuse methods that contain 49 pre-trained VLMs and 17 target datasets. Experiments demonstrate that the proposal can achieve state-of-the-art VLM selection performance, and the ability to deal with downstream tasks could grow with the scale of the model hub, showing the potential of building large model hubs with advanced model selection mechanisms.
@@
Code Availability Statement
The implementation code of benchmark and our proposal for the MLL paradigm is available at https://github.com/LAMDASZ-ML/MLL.
Acknowledgements
This research was supported by National Science Foundation of China (62306133, 62250069) and Key Program of Jiangsu Science Foundation (BK20243012). We would like to thank the reviewers for their constructive suggestions.
Impact Statement
This paper aims to explore how to select and reuse pre-trained VLMs for a specific downstream task. We are convinced that the majority of the impacts are positive, which is not necessary to specifically highlight any particular impacts in this paper. Moreover, we expect that the responsible use of technology will naturally promote discussions around best practices and regulations for implementing methods.